Find a degree 3 polynomial with coefficient of x^3 equal to 1 and zeros -3, −4i, and 4i. Simplify your answer so that it has only real numbers as coefficients.
Answer by stanbon(75887)

Find a degree 3 polynomial with coefficient of x^3 equal to 1 and zeros -3, −4i, and 4i. Simplify your answer so that it has only real numbers as coefficients.
---------
f(x) = (x+3)(x-4i)(x+4i)
---
f(x) = (x+3)(x^2+16)
-----
f9x) = x^3+3x^2+16x+48
=========================
Cheers,
Stan H.
=================